Drew burnt 1600 cal Friday playing one hour basketball and canoeing for two hours Saturday he spent two hours playing basketball in three hours canoeing in Byron 2700 cal how many calories did you burn per hour when playing basketball how many calories did you burn for our when canoeing canoeing


1
Expert's answer
2022-05-09T16:15:02-0400

Here we should create system:

where x - amount of calories burned playing basketball for 1 hour

y- amount of calories burned while canoeing for 1 hour

x+2y=1600

2x+3y=2700

express x through y

x=1600-2y and put that in the second equation

2(1600-2y)+3y=2700

3200-4y+3y=2700

y=500

x=1600-1000=600

x=600

y=500